# Population demographics of white dwarf binaries with intermediate separations: Gaia constraints on post-AGB mass transfer

### v1

@@ 3.1.1 Color excess distribution
Figure 5: Distribution of color excess for systems in the non-class I sample (Left) and those that end up in the final NCE sample (Right). We see that on average, triples have larger color excess and are therefore effectively excluded from the final sample. However, we also find that over half of true WD+MS binaries are removed in the process. Moreover, the distribution for the observed non-class I sample is significantly broader than that of our model, possibly due to inaccuracy in the measured metallicities (Section 3.1.1, Appendix A.3).
@@
Figure 4 shows the reported color excess of the observed non-class I sample as a function of $M_{1}$, [Fe/H], $P_{\rm orb}$, and eccentricity. Recall that Shahaf et al. (2024) set [Fe/H] = 0.0 for $\sim 20\%$ of sources with unreliable metallicites, causing the vertical streak at this value. Several other features stand out. Firstly, we observe a significant anti-correlation between color excess and [Fe/H]: essentially all sources with [Fe/H] $<-0.5$ have significant color excess. The source of this trend is uncertain, with one possibility being that the assumed relationship between color and metallicity is too strong in the PARSEC models (i.e. metal-poor stars are predicted to be bluer than they really are). Metal-poor stars are also older and thus on average are more evolved than they would be at the assumed age of 2 Gyr. This means that the observed sample is likely to be biased against metal-poor binaries. Second, we see there is a correlation between color excess and eccentricity. Hierarchical triples are expected to have higher eccentricities on average than WD+MS binaries, so this trend likely reflects an increasing fraction of triples at high color excess. A further discussion of this trend can be found in Appendix A.1. We also see that there is a cloud of sources with primary masses between $\sim 0.6-1.2\,M_{\odot}$ with systematically negative color excess, suggesting that the PARSEC models predict stars to be redder than they are in reality.
@@
Since we use PARSEC isochrones to calculate both the true and expected colors of the simulated binaries, the color excess distribution of our simulated WD+MS binaries peaks at zero, as expected. To properly reproduce the observed relation with primary mass would require identifying and correcting the underlying cause of the inaccurate colors predicted by the PARSEC models, which we do not attempt to do in this work. Instead, we apply an empirical correction to the color excess of our simulated systems. This is detailed in Appendix A.2. We note that the distributions of key parameters in the observed and modeled populations are comparable prior to the cut on color excess (i.e. the non-class I sample), so details of this correction do not play a critical role in our conclusions. However, the correction does affect the total number of systems that enter the final NCE sample and therefore adds uncertainty to the inferred occurrence rate of wide post-MT systems.
@@
The distributions of color excess for systems that enter the modeled (Section 5) and observed non-class I and final NCE samples are plotted in Figure 5. As expected, we see that on average, triples have significant non-zero color excess which results in the vast majority of them being excluded from the final sample. At the same time, around half of the simulated WD+MS binaries also have non-zero color excess and are excluded, which is primarily due to the fixed age assumption made in the calculation of color excess (Section 3.1) and suggests a less stringent cut may have been possible. In addition, the observed distribution is broader. One possible cause of this is that there is a discrepancy between the true and measured metallicities. This may also be responsible for the difference in the metallicity distributions of the observed and model samples (Figure 7). While there is no obvious cause of this, we note that the Zhang et al. (2023) metallicities were calculated assuming gaia_source parallaxes (as opposed to those from the NSS catalog) and did not account for the presence of light from luminous companions in the XP spectra. Since the uncertainties on these parallaxes are underestimated (El-Badry, 2025), it is likely that they metallicities are less accurate than their uncertainties suggest. The effect of introducing such a discrepancy is explored in Appendix A.3 but it does not significantly affect the distributions of the other parameters in the final NCE sample. However, a broader color excess distribution reduces the total number that enter the sample, adding an uncertainty of $\sim 5\%$ on the inferred fraction of post-AGB MT systems that end up in wide orbits, on top of the that coming from the choice of an empirical correction described above.
3.1.2 Deficit of triples in the simulated population
@@
Figure 6: AMRF against primary mass for the observed (left) and mock (fiducial model; right) non-class I samples. The points are colored by their color excess. The shaded region is the expected location for systems hosting dark secondaries with masses between $0.45-0.75\,M_{\odot}$ (i.e. WD+MS binaries). The dashed and dotted lines are upper limits to systems consistent with hosting a single MS secondary and an inner binary of two MS stars respectively. In both samples, there is a group of sources with near-zero color excess and located in the shaded region corresponding to WD+MS binaries. The others with large positive color excess values are most likely triples, which are rarer in our simulated population than in the observed sample.
@@
In Figure 6, we plot the AMRF against primary mass for sources in the observed and mock (Section 5) non-class I samples, colored by the color excess. The purple strip corresponds to the expected location of systems with dark $0.45-0.75\,M_{\odot}$ companions. The dash and dotted lines are the boundaries below which systems are consistent with hosting a single MS star and an inner binary respectively (i.e. non-class I systems lie above the dotted line; Section 3). These plots can be compared to Figure 5 in Shahaf et al. (2024).
@@
In both samples, there is a high density of sources occupying the strip with color excess close to zero, corresponding to the true WD+MS binaries. The distribution is slightly narrower and shifted to the left for the mock sample, which is reflected in the primary and WD mass distributions as discussed further in Section 6.
@@
Here, we focus on the sources with positive color excess values located outside of the strip and in between the two boundaries, which are presumably the triples. For the simulated sample, the vast majority of triples do indeed fall in this group (Figure 5). However, we see that there are significantly more such likely triples in the observed sample, likely reflecting an imperfect model of compact hierarchical triples in our simulated sample. We could increase the total number of these sources in our simulated population if we alter the mass ratio distribution from which we draw the outer tertiary’s mass so that there are fewer systems for which both stars in the inner binary have evolved into WDs (Section 2.5). Still, this would be unable to reproduce the pileup of observed sources at $M_{1}\sim 0.95\,M_{\odot}$, also seen in Figure 6 of Shahaf et al. (2024). This feature may suggest an additional structure in triple fraction or triple mass ratio distribution for compact triples that is not accounted for in our model. Alternatively, it could be the result of overestimated primary masses for the observed sources in the binary_masses catalog. Since almost all triples are effectively removed by the cut on color excess, these discrepancies have a minimal effect on our final sample but future work to study the nature of these sources may be informative.
@@ 4 Results of E24 model
We first consider the original mock NSS catalog from E24, with the addition of hierarchical triples (Section 2.5). This will be referred to as the E24 model. In Figure 7, we compare distributions of several parameters for systems in the observed (filled, light blue) and the simulated E24 (red dashed border) NCE samples. We also plot the results of our new “fiducial” model (purple solid border) where several improvements have been made on the E24 model, as described in Section 5. $M_{\rm 2,dark}$ is the secondary mass implied by the astrometric solution assuming a completely dark companion. In these plots, the total number of systems has been adjusted to be equal for easier comparison of the shapes of distributions. Here, the total number of systems in the mock sample is within $8\%$ of the true sample, suggesting that the assumed $10\%$ fraction of post-AGB MT systems in wide orbits is reasonable, given the other assumptions of the model (see discussion towards the end of Section 6.1).
We see that there is a deficit of systems with $P_{\rm orb}<300\,$d and $e>0.2$ in the E24 model. The color excess cut removes the vast majority of hierarchical triples that enter the non-class I sample so that they make up less than 0.5% of the NCE sample. Therefore, we can conclude that the deficit of systems with $e>0.2$ cannot be attributed to contaminating triples. The distribution of color excess is discussed further for our new fiducial model in Section 6.1 (Figure 5). The lack of shorter period systems can be addressed by modifying the conditions that lead to the formation of wide post-MT systems, as will be described in Section 5.1 and illustrated in Figure 8.
The model significantly overpredicts the fraction of WDs with masses $\gtrsim 0.8\,M_{\odot}$, which we address for our fiducial model in the following sections. This deficit of massive WDs in the AMRF sample compared to the field WD population has been explored recently by Hallakoun et al. (2024) (Section 6.1). Massive WDs are over-represented in the AMRF sample as they produce larger orbits and can therefore be identified as WDs even when they orbit more massive/luminous stars, so the dearth of massive WDs is highly significant.
@@
We also see that there is a slight discrepancy between the observed and simulated [Fe/H] distributions. This may be attributed to uncertainties in the underlying stellar population simulated by Galaxia or those in the measured metallicites reported by Zhang et al. (2023) (Section 6.1).
Figure 7: Distributions of stellar and orbital parameters in the real NCE sample from Shahaf et al. (2024), and mock samples obtained from processing the original mock NSS catalog from E24 (with the inclusion of hierarchical triples; Section 2.5), and our updated fiducial model (Section 5) Overall, the fiducial model improves on the major discrepancies present between the distributions of the E24 model and observed sample.
@@ 5 Modifications to the E24 model
The key components of our updated model are summarized in Figure 8, which we describe in the following.
5.1 Orbital evolution under MT
Figure 8: A schematic summarizing the assumed orbital evolution of the zero-age binary population in our fiducial model which lead to the formation short-period and wide post-MT binaries. Detailed descriptions of the components can be found in Section 5.
@@ 5.1 Orbital evolution under MT
Figure 9: Maximum radius that a star reaches on the RGB and AGB as a function of its initial mass. In our model, systems with initial pericenter separations below the RGB curve are assumed to become short-period PCEBs, while a fraction of those in between the two curves end up as wide PCEBs.
@@
As mentioned in Section 2.3, the outcome of binary mass transfer is expected to depend on the structure of the donor’s envelope. In particular, compared to red giants, AGB stars can have very loosely bound envelopes that may be ejected with very little energy (e.g. Belloni et al., 2024; Yamaguchi et al., 2024a). We therefore identify and separate the orbital evolution of binaries which interact when the donor is on the RGB versus the AGB by comparing their initial separations to the maximum stellar radii at each of these phases.
@@
We use the standard set of MIST evolutionary tracks (with rotation; Dotter 2016; Choi et al. 2016) to predict the maximum radius reached on the RGB and AGB, $R_{\rm RGB,max}$ and $R_{\rm AGB,max}$ respectively, for a given initial stellar mass (i.e. mass of the WD progenitor). This is plotted in Figure 9. We fit two cubic spline functions for the RGB track above and below $2\,M_{\odot}$. For the AGB track, there is large scatter due to thermal pulses, particularly above $\sim 5\,M_{\odot}$. To avoid over-fitting the data, we excluded local dips in the maximum radii at these masses while fitting a single cubic spline. We note that progenitors with masses above $\sim 5\,M_{\odot}$ make up a small fraction of the total population ($\sim 3\%$) and thus the effect of this choice on the statistics of the final population is small.
We modify the conditions that lead to the formation of short ($P_{\rm orb}=1\,d$) and longer period post-interaction binaries from those described in Section 2. Firstly, instead of using a simple cut on the semi-major axis, we consider the Roche lobe radius at periastron, $r_{\rm L,peri}=f_{q}a_{i}(1-e)$ (where $f_{q}$ is the Eggleton factor; Eggleton 1983). This is reasonable as interaction most likely begins when the two components are closest to each other. We assume that systems which first interact when the WD progenitor is on the first giant branch, where $r_{\rm L,peri}<R_{\rm RGB,max}$, end up with short (111\,1d) periods. Meanwhile, a fraction of systems that first interact on the AGB, where $R_{\rm RGB,max}<r_{\rm L,peri}<R_{\rm AGB,max}$, may form wide post-interaction binaries. In the E24 model, 10% of such systems were randomly chosen to end up in orbits with final separations that are a constant fraction, $\beta=0.5$, of their initial separations. In our updated fiducial model, we instead assume that all systems that interact with an AGB donor evolve with a $\beta$ that depends on the initial mass ratio, $q_{\rm init}=M_{\rm 2,init}/M_{\rm 1,init}=M_{\rm 1,final}/M_{\rm WD,prog}$ such that it increases with higher $q_{\rm init}$. This means that on average, systems with more massive accretors/less massive donors experience less orbital shrinkage and form wider post-MT systems, which are more likely to receive astrometric orbits. This choice is motivated by the discrepancy in the component mass distributions of the final WD+MS binaries in the observed sample and the E24 model (Figure 7). We use a sigmoid function, plotted in Figure 10, which provides a steep dependence on $q_{\rm init}$ required to match the observed deficit of low mass primaries and high-mass WDs. Overall, this results in $13\%$ of post-AGB interaction systems forming “wide” orbits with $P_{\rm orb}=100-1000\,$d. Although a larger fraction of these systems have $q_{\rm init}>0.38$ above which $\beta$ is large (approaching 0.63), the majority end up in periods $>1000\,$d due to the decrease in mass resulting from the formation of the WD. While we ultimately chose a steeply rising sigmoid function that is step-like, in Appendix B, we show the effects of varying the form of this function on the final results.
While it is possible to achieve a similar effect by preserving the E24 setup (leaving $\beta$ constant, instead allowing the fraction that form wide orbits to vary with $q_{\rm init}$), our new choice is more physically motivated. For example, the $\alpha$-formalism prescription for the common envelope phase, which is an argument of energy conservation, predicts a final separation that depends on the component masses and thus implicitly on their ratio. In Section 5.3, we describe an alternative model for $\beta$ built using on the $\alpha$-formalism.
Furthermore, the mass ratio at which the step occurs ($q_{\rm init}=0.38$ in our fiducial model) can be understood as the critical mass ratio (e.g. Ge et al., 2020; Temmink et al., 2023) above which mass transfer remains stable and orbits stay wide, and below which CEE ensues and orbits shrink significantly. As we will discuss further, it is debatable whether stable mass transfer or CEE is a more accurate description of the binary interaction that leads to the formation of wide post-MT binaries. The key input for our model is simply that binaries with extreme mass ratios shrink to small orbits, whereas those with closer-to-equal mass ratios remain relatively wide.
Figure 10: The ratio $\beta$ of the final to initial separation as a function of the initial mass ratio $q_{\rm init}=M_{\rm 2,init}/M_{\rm 1,init}$. The curve is a Sigmoid function whose functional form is given at the bottom right corner of the figure.

### v3

@@ 3.1.1 Color excess distribution
Figure 5: Distribution of color excess for systems in the non-class I sample (Left) and those that end up in the final NCE sample (Right). On average, triples have larger color excess and are therefore effectively excluded from the final sample. However, we also find that more than half of true WD+MS binaries are removed in the process. The distribution for the observed non-class I sample is significantly broader than that of our model, possibly due to inaccuracy in the measured metallicities (Section 3.1.1 and Appendix A.3).
@@
Figure 4 shows the reported color excess of the observed non-class I sample as a function of $M_{1}$, [Fe/H], $P_{\rm orb}$, and eccentricity. Recall that Shahaf et al. (2024) set [Fe/H] = 0.0 for $\sim 20\%$ of sources with unreliable metallicites, causing a vertical streak at [Fe/H] = 0.0. Several other features stand out. Firstly, we observe a significant anti-correlation between color excess and [Fe/H]: essentially all sources with [Fe/H] <0.5<-0.5 have significant color excess. The source of this trend is uncertain, with one possibility being that the assumed relationship between color and metallicity is too strong in the PARSEC models (i.e. metal-poor stars are predicted to be bluer than they really are). Metal-poor stars are old, and thus on average are more evolved than they would be at the assumed age of 2 Gyr. This means that their color excess will be overestimated on average, and that the observed sample is likely to be biased against metal-poor WD+MS binaries.
Second, there is a correlation between color excess and eccentricity. Hierarchical triples are expected to have higher eccentricities on average than WD+MS binaries, so this trend likely reflects an increasing fraction of triples at high color excess. Further exploration of the trend can be found in Appendix A.1.
We also observe a cloud of sources with primary masses of $0.6-1.2\,M_{\odot}$ with systematically negative color excess, suggesting that the PARSEC models predict stars to be redder than they are in reality.
@@
Since we use PARSEC isochrones to calculate both the true and expected colors of the simulated binaries, the color excess distribution of our simulated WD+MS binaries peaks at zero, as expected. To fully reproduce the observed relation between color excess and primary mass would require identifying and correcting the underlying cause of the inaccurate colors predicted by the models, which we do not attempt to do in this work. Instead, we apply an empirical correction to the color excess of our simulated systems. This is detailed in Appendix A.2. We note that the distributions of key parameters in the observed and modeled populations are comparable prior to the cut on color excess (i.e. the non-class I sample), so details of this correction do not play a critical role in our conclusions. However, the correction does affect the total number of systems that enter the final sample and therefore adds uncertainty to the inferred occurrence rate of wide post-MT systems.
The distributions of color excess for systems that enter the simulated (Section 5.4.4) and observed non-class I and final NCE samples are plotted in Figure 5. On average, triples have significant non-zero color excess, which results in the vast majority of them being excluded from the final sample. At the same time, around half of the simulated WD+MS binaries also have non-zero color excess and are excluded, which is primarily due to the fixed age assumption made in the calculation of color excess (Section 3.1) and suggests a less stringent cut may have been possible. In addition, the observed distribution is broader.
@@
A possible explanation is that there is a discrepancy between the true and measured metallicities. This may also be responsible for the difference in the metallicity distributions of the observed and model samples (see discussion around Figure 7). We note that the Zhang et al. (2023) metallicities were calculated assuming gaia_source parallaxes (as opposed to those from the NSS catalog) and did not account for the presence of light from luminous companions in the XP spectra. Since the uncertainties on these parallaxes are underestimated (El-Badry, 2025), it is likely that the metallicities are less accurate than their uncertainties suggest. The effect of introducing such a discrepancy is explored in Appendix A.3; it does not significantly affect the distributions of the other parameters in the final NCE sample. However, a broader color excess distribution reduces the total number of systems that enter the final sample, adding another source of uncertainty on the inferred fraction of post-AGB MT systems that end up in wide orbits.
3.1.2 AMRF distributions
@@
Figure 6: AMRF against primary mass for the observed (left) and mock (empirical model; right) non-class I samples. The points are colored by their color excess. The shaded region is the expected location for systems hosting dark secondaries with masses between $0.45-0.75\,M_{\odot}$ (i.e. WD+MS binaries). The dashed and dotted lines are upper limits to systems consistent with hosting a single MS secondary and an inner binary of two MS stars, respectively. In both samples, there is a group of sources with near-zero color excess located in the shaded region, corresponding to WD+MS binaries. Sources with large positive color excess values are most likely triples, which are rarer in our simulated population than in the observed sample.
@@
In Figure 6, we show the AMRF and primary mass for sources in the observed and mock (Section 5) non-class I samples, colored by color excess. The shaded strip corresponds to the expected location of systems with dark $0.45-0.75\,M_{\odot}$ companions. The dashed and dotted lines are the boundaries below which systems are consistent with hosting a single MS star and an inner binary, respectively (i.e. non-class I systems lie above the dotted line). These plots can be compared to Figure 5 of Shahaf et al. (2024).
@@
In both panels, there is a high density of sources occupying the shaded strip with color excess close to zero, corresponding to true WD+MS binaries. The distribution in the mock sample is somewhat narrower and has a shallower slope than in the observed sample. This tension reflects a stronger correlation between primary and secondary masses in the mock sample than in the observed sample, as we discuss further in Section 6.4.
Here, we focus on the sources with positive color excess values located outside of the strip and in between the two boundaries, which are presumably the triples. For the simulated sample, the vast majority of triples do indeed fall in this clump. However, there are significantly more such likely triples in the observed sample, likely reflecting an imperfect model of compact hierarchical triples in our simulated sample.
@@
We could increase the total number of these sources in our simulated population by altering the mass ratio distribution from which we draw the outer tertiary’s mass, resulting in fewer systems for which both stars in the inner binary have evolved into WDs (Section 2.5). Still, this would fail to reproduce the pileup of observed sources at $M_{1}\sim 0.95\,M_{\odot}$, also seen in Figure 6 of Shahaf et al. (2024). This feature may reflect additional structure in compact triple population that is not accounted for in our model. Alternatively, it could be the result of biased primary masses for the observed sources in the binary_masses catalog. Since almost all triples are effectively removed by the cut on color excess, these discrepancies have a minimal effect on our final sample.
@@ 4 Results of E24 model
We first consider the original mock NSS catalog from E24, with the addition of hierarchical triples (Section 2.5). This will be referred to as the “E24 model”. In Figure 7, we plot the distributions of several stellar and orbital parameters of the NCE sample resulting from this model (red dashed border) on top of those observed (light blue, filled). In Figure 8, we compare these samples (panels (A) and (B)) in the $P_{\rm orb}-M_{\rm 2,dark}$ plane, where $M_{\rm 2,dark}$ is the secondary mass implied by the astrometric solution assuming a completely dark companion (i.e. the approximate WD mass). The distribution of each parameter is plotted on the corresponding axis, and points are colored by companion mass.
The total number of systems in the mock sample is within $8\%$ of the true sample, suggesting that the assumed $10\%$ fraction of post-AGB MT systems in wide orbits is reasonable, given the other assumptions of the model (see discussion towards the end of Section 3.1.1).
We see that there is a deficit of systems with $P_{\rm orb}<300\,$d and $e>0.2$ in the E24 model. The color excess cut removes the vast majority of hierarchical triples that enter the non-class I sample so that they make up less than 0.5% of the NCE sample. Therefore, we can conclude that the deficit of systems with $e>0.2$ cannot be attributed to contaminating triples. The distribution of color excess is discussed further for our udpated empirical model in Section 3.1.1 (Figure 5). The lack of shorter period systems can be addressed by modifying the conditions that lead to the formation of wide post-MT systems, as will be described in Section 5.4.
The model significantly over-predicts the fraction of WDs with masses $\gtrsim 0.8\,M_{\odot}$, which we address in our updated models as described in Section 5. This deficit of massive WDs in the AMRF sample compared to the field WD population has been explored recently by Hallakoun et al. (2024). Massive WDs are over-represented in the AMRF sample as they produce larger orbits and can therefore be identified as WDs even when they orbit more massive/luminous stars, so the dearth of massive WDs is highly significant.
@@
We also see that there is a slight discrepancy between the observed and simulated [Fe/H] distributions. This may be attributed to uncertainties in the underlying stellar population simulated by Galaxia or those in the measured metallicites reported by Zhang et al. (2023) (Section 3.1.1).
Figure 7: Distributions of stellar and orbital parameters in the observed NCE sample from Shahaf et al. (2024), and mock samples obtained from processing the original mock NSS catalog from E24 (with the inclusion of hierarchical triples; Section 2.5), the stable MT model (Section 5.4.1), and our empirical model; Section 5.4.4). For comparison purposes, the total number of systems in all samples are made to be equal. Overall, the empirical model best reproduces the features of the observed sample.
Figure 8: Orbital period and WD mass, colored by companion mass, for systems in the observed sample (Panel A; Shahaf et al. 2024), the mock sample resulting from the E24 model (Panel B; Section 4), and various other models of the orbital evolution tested in this work (Panels C - F; Section 5.4). For ease of comparison, the total number of systems in all samples are adjusted to be equal, and the observed sample is shown in gray across all panels.
@@ 5 Modifications to the E24 model
We tested several modifications to the E24 model. We first describe the shared changes made across all of our new models.
5.1 Interaction on the RGB vs. AGB
@@
Figure 9: Maximum radius that a star reaches on the RGB and AGB as a function of its initial mass. The points are obtained from MIST evolutionary tracks, and the curves are splines fitted to the points used for interpolation. In our model, systems with initial pericenter separations below the RGB curve are assumed to become short-period PCEBs, while a fraction of those in between the two curves end up as wide PCEBs.
@@
As mentioned in Section 2.3, the outcome of binary mass transfer is expected to depend on the structure of the donor’s envelope. In particular, compared to red giants, AGB stars can have very loosely bound envelopes that may be ejected with very little energy (e.g. Belloni et al., 2024; Yamaguchi et al., 2024a). AGB donors may also be able to maintain stability over more extreme mass ratios (e.g. Temmink et al., 2023). We therefore identify and separate the orbital evolution of binaries which interact when the donor is on the RGB versus the AGB by comparing their initial separations to the maximum stellar radii at each of these phases.
@@
We use the standard set of MIST evolutionary tracks (with rotation; Dotter 2016; Choi et al. 2016) to predict the maximum radius reached on the RGB and AGB, $R_{\rm RGB,max}$ and $R_{\rm AGB,max}$ respectively, for a given initial stellar mass (i.e. mass of the WD progenitor). This is plotted in Figure 9. We fit two cubic spline functions for the RGB track above and below $2\,M_{\odot}$. For the AGB track, there is large scatter due to thermal pulses, particularly above $\sim 5\,M_{\odot}$. To avoid such features, we fit the maximum value in a sliding window spanning five data points. We note that progenitors with masses above $5\,M_{\odot}$ make up a small fraction of the total population ($\sim 3\%$) and thus the effect of this choice on the statistics of the final population is small.
Furthermore, we consider the Roche lobe radius at periastron, $r_{\rm L,peri}=f_{q}a_{i}(1-e)$ (where $f_{q}$ is the Eggleton factor; Eggleton 1983, $a_{i}$ is the initial semi=major axis, and $e$ is the eccentricity). This is reasonable as interaction most likely begins when the two components are closest to each other. We assume that systems with $r_{\rm L,peri}<R_{\rm RGB,max}$ first interact when the WD progenitor is on the red giant branch. Meanwhile, systems with $R_{\rm RGB,max}<r_{\rm L,peri}<R_{\rm AGB,max}$ first interact with an AGB donor. In the E24 model, 10% of the systems in the latter group were randomly chosen to end up in orbits with final separations that are a constant fraction, $f_{\rm shrink}=0.5$, of their initial separations. We modify this evolution, testing several different models, which are described in Section 5.4.
5.2 Donor mass at the onset of MT
Figure 10: Stellar mass as a function of initial mass and radius on the AGB. The markers are from MIST evolutionary tracks, while the color scale shows our interpolation. We see significant mass loss occurring as the star expands to large radii near the end of the AGB phase.
The models of orbital evolution described in Section 5.4 all depend on the donor mass, or the system mass ratio, at the onset of MT. These quantities are generally different from their initial values, as the donor may have experienced significant mass loss via winds. We determine the donor mass at Roche lobe overflow, $M_{\rm 1,MT}$, by interpolating MIST tracks to predict the stellar mass given the initial mass and radius at the time of Roche lobe overflow (Figure 10). For each system which interacts on the RGB or AGB, we use 2D interpolation with the scipy.interpolate.RBFInterpolator package to predict the donor’s mass at the time of Roche lobe overflow. Figure 10 shows the interpolated function on the AGB. During thermal pulses, the mass-radius relation for a given initial mass can be non-monotonic. However, this is not problematic for our models, as we only consider interactions where the radius is larger than at any earlier point in the evolution (otherwise interaction would have occurred earlier).
5.3 Eccentricity evolution